OpenStudy (anonymous):

Can someone please show me how to prove that the following integral = 1

OpenStudy (anonymous):

\[\theta^k (-y^-k), 0<=y<=2\]

OpenStudy (dumbcow):

?? i don;t see an integral

OpenStudy (anonymous):

??? Where is the integral ??

OpenStudy (anonymous):

Sorry - I'll attach what I started with and my workings. I am confused about where to apply the limits (and even if the limits are correct)

OpenStudy (anonymous):

Can you only attach the question separately because I am unable to distinguish what you have worked out and what is the question given

OpenStudy (anonymous):

Sure no problem

OpenStudy (anonymous):

k and theta are constants,Am I right ?

OpenStudy (anonymous):

That is my understanding, yes

OpenStudy (anonymous):

First of all I would like to say that you have gone wrong in your integration. \[k \theta^{k} \int\limits_{}^{}y^{-k-1} dy= k \theta^{k}{{ y^{-k-1+1} }\over {-k+1-1}} = -\theta^{k}y^{-k}\]

OpenStudy (dumbcow):

ahh you beat me to it...and the limits are from theta to infinity

OpenStudy (anonymous):

Sorry @SKMC you have done it correctly @dumbcow , thanks for telling the limits.

OpenStudy (anonymous):

@dumbcow , Can you explain how you have arrived at these limits ??

OpenStudy (dumbcow):

@shivam_bhalla , given theta >0 and y>theta so min_y must be theta and max_y has no limit

OpenStudy (anonymous):

but you have not taken into account that for theta>0 and k>2 --> f(y)=0

OpenStudy (anonymous):

This is the point where i became unsure how to proceed

OpenStudy (dumbcow):

oh i see, maybe i misread that part hmm but if you assume f(y)>0 when y>theta it should still work and the integral evaluates to 1 with those limits

OpenStudy (anonymous):

@dumbcow, you are right. The integral evaluates to 1

OpenStudy (anonymous):

I dont understand how that evaluates to 1

OpenStudy (dumbcow):

\[k \theta^{k}\int\limits_{\theta}^{\infty}y^{-k-1} dy = k \theta^{k} |_\theta^{\infty} \frac{y^{-k}}{-k} = k \theta^{k}(0-\frac{\theta^{-k}}{-k}) = \theta^{k-k} = 1\]

OpenStudy (anonymous):

So from my workings, where I got \[\theta^k(-y^-k)\] my next line should have been \[\theta ^{k-k}\] instead of \[-(\theta/y)^{k}\]and from \[\theta ^{k-k}\] \[\theta =1?\]

OpenStudy (dumbcow):

anything to zero power is 1

OpenStudy (anonymous):

Many Thanks for your help shivam_bhalla and dumbcow- i appreciate your expertise and your patience :)